OpenStudy (anonymous):

An arithmetic sequence is represented by the explicit formula A(n) = 2 + 9(n - 1). What is the recursive formula?

OpenStudy (anonymous):

A(n+1)= 2+ 9((n+1)-1) - (2 + 9(n-1)) =9 i.e is A(n+1)=A(n) + 9;

OpenStudy (anonymous):

thanks